Home‎ > ‎Crew 153‎ > ‎


Commander's Report
Date: April 16, 2015
Crew Commander Name: Bastien MATHURIN

Hello Earth,

We had a great cold day on Mars. Today Gaspard and Martin are going to cook pasta with Bolognese sauce. I hope that it won't be as spicy as yesterday's meal.

We had requested an EVA for today. Unfortunately the weather was not on our side and we had to change the objectives of this EVA. We requested the use of ATVs in order to go in an old streambed located about 5 miles away from the MDRS's location in order to perform hydraulic measurements for my experiment. However it rained and snowed all the day leading to impassable roads for safety concerns. We've finally chosen to go outside just in front of the hab for Auriane and Gaspard experiments. We went out only for one hour instead of two. They both got some good results for their respective scientific experiments.

Regarding the other experiments, it seems that all is going well. Florian succeeded to make bacteria found on Martian rocks grow. He is currently trying to isolate them for a future characterization. He is also testing all his staining protocols. Auriane will begin to develop her location tool without using a GPS. She will be able (we hope) to map the path used during an eva - but the localization won't be as good as it is using a GPS. Concerning Martin's experiment, all his devices are ready, but the weather is not good enough to make measurements using the telescope. He just can't get a "ray of black night". Unfortunately, since the accuracy of the spectroscope is limited, he will first try to take spectra of planets and stars. If he has enough time, he will try to measure galaxies redshift at the end of our mission. Romain is currently writing his video novel despite his illness. It does not progress as quickly as desired but at least it goes ahead. Gaspard has improved his wifi intensity mapping device and he is going to analyze the results obtained during today's EVA as soon as possible. As far as I'm concerned, my hydraulic experiment doesn't seems to be leading to satisfying results: the measurements of the geometry are so inaccurate that nothing can be concluded for the moment. I will try again during the next days.

Following our visit to the clinic yesterday for Romain's illness, it seems that he is already feeling a bit better. He still has a sore throat, but despite this problem, there seems to be a positive evolution: he already feels stronger than before.

Concerning the issue about the water supply in the Hab, there is no news since yesterday. We're supposed to receive a new pump in the days to come. Our Crew Engineer feels comfortable enough to replace the pump when it will be delivered. A new pump will be much more convenient for our life in the station.

On the other hand, as I always mentioned, the crew is still in high spirits!

Ad astra!